# BiMark: Unbiased Multilayer Watermarking for Large Language Models

### v1

@@ 1 Introduction
Large Language Models (LLMs) (OpenAI, 2022; Team, 2024b) have recently emerged in advancing cutting-edge technologies, and their rapid evolution has made LLM-generated content increasingly indistinguishable from human-created text. However, pressing security challenges (Zhang et al., 2023b; Wei et al., 2025; Gong et al., 2025; Luo et al., 2024; Zhang et al., 2025a, b, 2023a, 2024b; Pan et al., 2025) have been raised with the rapid progression of AI technologies (Zheng et al., 2024b; Zhang et al., 2024a; Zheng et al., 2024a, 2023, 2025; Guan et al., 2023), such as the misuse of LLMs (Tang et al., 2024) and synthetic data pollution (Shumailov et al., 2024). Therefore, recent legislation, such as the AI Act of the European Union (European Commission, 2023) and the executive order of the U.S. Department of Commerce (The White House, 2023), mandates the implementation of technical measures to mark and detect LLM-generated content. Thus, it appeals to LLM watermarking (Zhang et al., 2024e; Tang et al., 2024), a promising method to address these challenges, offering a proactive approach to embedding identifiable patterns in LLM-generated text (Liu et al., 2024).
@@
Inference- Time / Aaronson (2023)
@@
Cumulative Probability $\gamma$-reweight (Hu et al., 2023)
DiPmark (Wu et al., 2023)
Token Counting GINSEW (Zhao et al., 2023)
@@
Depending on whether it is integrated with the language model, watermarking can be categorized into inference-time watermarking and post-hoc watermarking (Tang et al., 2024). Compared to post-hoc methods that require an additional auxiliary module (Abdelnabi & Fritz, 2021; Yang et al., 2022; Yoo et al., 2023a), inference-time watermarking (Dathathri et al., 2024; Aaronson, 2023; Kirchenbauer et al., 2023) can embed detectable patterns linked to a secret key by modifying the LLM token sampling process in a plug-and-play manner, thereby efficiently achieving a cryptographically guaranteed detectable secret watermark (Aaronson, 2023; Kirchenbauer et al., 2023).
In general, implementing language model watermarking necessitates three key requirements: 1) Text quality preservation: The integration of watermarking mechanisms must maintain the fundamental utility and performance of language models (Hu et al., 2023; Kuditipudi et al., 2023). 2) Model-agnostic detection: The watermark verification process should function independently of the generating model’s architecture, parameters, or access rights (Kirchenbauer et al., 2023; Kuditipudi et al., 2023). 3) Message embedding capacity: The watermarking scheme should accommodate sufficient capacity to encode crucial metadata such as model identity, generation timestamp, and content provenance (Bob & Dan, 2024; Pang et al., 2024).
Note that it is not trivial to satisfy the above three requirements, as there is a trade-off between text quality preservation and message embedding capacity (Yoo et al., 2023b; Qu et al., 2024; Kirchenbauer et al., 2023). Some existing works (Yoo et al., 2023b; Qu et al., 2024) achieve efficient multi-bit watermarking by embedding one message bit per token through hashing. However, they rely on text quality-compromising zero-bit watermarking (Kirchenbauer et al., 2023) as a foundation, leading their enhancement of message embedding capacity inevitably come at the cost of text quality preservation. Note that, although other methods (Fernandez et al., 2023; Fairoze et al., 2023; Wang et al., 2023) can mitigate this issue to some extent by associating messages with secret keys or hash functions, they lack detection efficiency due to using messages non-invertibly.
@@ 2 Related Work
LLM watermarking aims to address the urgent need to differentiate model-generated text from human-written content. Text watermarking (Fang et al., 2017; Ziegler et al., 2019; He et al., 2022b; Fairoze et al., 2023; Fridrich et al., 2004; Christ et al., 2024; Zhao et al., 2024) has evolved significantly since its early application in content protection. Here we review related work on inference-time watermarking around the three key capacities in LLM watermarking.
@@
Text quality preservation. Early watermarking methods such as Soft Red List (Kirchenbauer et al., 2023) significantly influenced the field but faced challenges with text quality preservation due to their biased nature. Recent work has addressed this challenge through two main approaches: unbiased reweighting (Hu et al., 2023; Wu et al., 2023) and distortion-free sampling (Kuditipudi et al., 2023; Zhao et al., 2024; Dathathri et al., 2024). Unbiased reweighting methods, such as $\gamma$-reweight (Hu et al., 2023) and DiPmark (Wu et al., 2023), modify the probability distributions of generated text to inject watermarks while ensuring that the expected probability distribution remains unchanged. Distortion-free sampling methods, such as Gumbel sampling (Aaronson, 2023; Kuditipudi et al., 2023) and SynthID (Dathathri et al., 2024), employ secret keys to guide the sampling process rather than modifying the probability distribution. Many of them enable model-agnostic detection, but lack multi-bit functionality.
@@
Model-agnostic detection. The ability to detect watermarks without accessing the original or an auxiliary model is crucial for practical deployment. Existing watermarking methods developed from unbiased reweighting (Hu et al., 2023; Wu et al., 2023) utilize cumulative probability to partition the vocabulary, which makes them unsuitable for model-agnostic detection. In the scenario of multi-bit watermarking, a similar concept critical for practical deployment is message-agnostic detection. That is, the ability to extract messages without access to the message space. Among existing multi-bit watermarking methods, only the works of Yoo et al. (2023b) and (Qu et al., 2024) satisfy this property through extracting messages from a voting matrix bit by bit. However, they are developed based on Soft Red List and thus without guaranteeing text quality.
@@
Message embedding capacity. Message embedding capacity enables information to be embedded in and extracted from watermarked text, which is essential for tracing text provenance (Cohen et al., 2025; Pang et al., 2024). Recent studies develop multi-bit watermarking for message embedding capacity by associating each message with a unique secret key (Fernandez et al., 2023), or incorporating messages into hash functions as inputs (Fairoze et al., 2023; Wang et al., 2023). These methods require access to message space for extraction, as the messages themselves are entangled with hash functions in a non-invertible manner.
@@
Yoo et al. (2023b) proposed a position allocation technique incorporating Soft Red List (Kirchenbauer et al., 2023), which can solve this problem. They allocate each token to a subunit of a message for watermark embedding, and extract the message bit by bit, which is in a message-agnostic way.  Qu et al. (2024) further incorporates error correction codes to enhance multi-bit watermarking resilience. However, these methods were designed using Soft-List, leading to a lack of unbiasedness. As shown in the Tab. 1, a comprehensive comparison between BiMark and related studies is presented, highlighting our contributions in this field.
@@ 3.1 Watermarking for LLMs
Watermarking LLM-generated text. To reliably distinguish between LLM-generated and human-written text, watermarking methods actively inject detectable signals using secret keys (Kirchenbauer et al., 2023; Aaronson, 2023) into LLM-generated text. During detection, through measuring the similarity between extracted signals and injected signals with the same keys, the watermark can be verified. Early works (Kirchenbauer et al., 2023; Kuditipudi et al., 2023; Hu et al., 2023) only focus on addressing the binary question of whether a text contains watermarks, referred to zero-bit watermarking. For more reliable text authenticity, watermarking methods that can carry informative messages are proposed, referred to as multi-bit watermarking  (Yoo et al., 2023b; Fairoze et al., 2023; Cohen et al., 2025).
@@ 3.2 Soft Red List Watermarking
If the z-score surpasses a predefined threshold, it suggests that the text is watermarked, since the proportion of green tokens deviates significantly from the expected value $\gamma$. When using the z-score for watermark detection, the corresponding p-value reflects the false positive rate (Type-I error), whereas the false negative rate (Type-II error) is influenced by the inherent entropy characteristics of the LLM (Kirchenbauer et al., 2023; Kuditipudi et al., 2023).
@@ 3.3 Watermarking via Unbiased Reweighting
Definition 3.1 ($n$-shot undetectable (Hu et al., 2023)).
@@
(Unbiased reweighting function (Hu et al., 2023)). Given a random variable $E$ and a reweighting function $R_{E}:\Delta_{\mathcal{V}}\rightarrow\Delta_{\mathcal{V}}$, we say that $R_{E}$ is an unbiased reweighting function if and only if for all probability distributions $P\in\Delta_{\mathcal{V}},~{}\mathbb{E}_{E}[R_{E}(P)]=P$.
@@ 3.4 Multi-bit Watermarking via Position Allocation
Many prior works (Fernandez et al., 2023; Fairoze et al., 2023; Wang et al., 2023; Yoo et al., 2023b; Qu et al., 2024) have explored multi-bit watermarking that extends beyond zero-bit detection. Among them, Yoo et al. (2023b) proposed Multi-bit Watermarking via Position Allocation (MPAC), which enables message embedding without additional latency compared to Soft Red List and allows message extraction in a message-agnostic manner. We examine this technique, as it inspired BiMark, which integrates it with XOR operations to achieve unbiased watermarking.
@@ 4.1 Bit-Flip Unbiased Reweighting
An unbiased reweighting function is the core component of unbiased watermarking. While unbiased reweighting helps preserve text quality in watermarking, existing methods  (Hu et al., 2023; Wu et al., 2023) based on cumulative probability distributions require access to model probabilities, making them unsuitable for model-agnostic detection.
@@
Formally, let a random variable ${e}$ represent a fair coin flip, where ${e}=1$ indicates heads and ${e}=0$ indicates tails. Given a vocabulary bipartition $(\mathcal{V}_{0},\mathcal{V}_{1})$, the bit-flip reweighting function Fig. 2 adjusts an original probability as follows:
$R_{\theta,e}(P_{M})(x)$ =$=\begin{cases}(1+\delta_{1})P_{M}(x)&\text{if }e=1\wedge x\in\mathcal{V}_{1},\\ (1-\delta_{0})P_{M}(x)&\text{if }e=1\wedge x\in\mathcal{V}_{0},\\ (1-\delta_{1})P_{M}(x)&\text{if }e=0\wedge x\in\mathcal{V}_{1},\\ (1+\delta_{0})P_{M}(x)&\text{if }e=0\wedge x\in\mathcal{V}_{0}.\end{cases}$ (1)
@@
For practical implementation, we introduce a base scaling factor $\tilde{\delta}\in[0,1]$ and derive $\delta_{1}$ adaptively based on $\tau$ from Eq. (3). However, two scenarios [^2] For ease of presentation, we assume $(1+\delta_{1})\tau>(1+\delta_{0})(1--\tau)$ and focus on the edge outlier caused by expanding $\mathcal{V}_{1}$. require special attention: (1) If $\tau=0$, we set $\delta_{1}=0$ since all tokens in $\mathcal{V}_{1}$ have zero probability and any non-zero $\delta_{1}$ would have no effect on the probability distribution and will break the symmetry of our reweighting scheme when ${e}=1$. (2) If $(1+\tilde{\delta})\tau>1$, we set $\delta_{1}=(1-\tau)/\tau$ which makes the probability of $\mathcal{V}_{1}$ become $1$ and the probability of $\mathcal{V}_{0}$ become 00, ensuring the reweighted distribution remains a valid probability distribution over two partitions. Consequently, the scaling factor $\delta_{1}$ becomes:
$\delta_{1}=\begin{cases}0&\text{if $\tau=0$},\\ (1-\tau)/\tau&\text{if }(1+\tilde{\delta})\tau>1,\\ \tilde{\delta}&\text{otherwise}.\end{cases}$ (3)
@@
For any probability distribution over a vocabulary $\mathcal{V}$ and a reweighting function $R_{\theta,e}$ defined above, we have $\mathbb{E}_{e}[R_{\theta,{e}}(P)]=P$.
@@ 4.2 Multilayer Unbiased Reweighting
Formally, given a sequence of independent vocabulary bipartitions $[(\mathcal{V}_{0}^{1},\mathcal{V}_{1}^{1}),\cdots,(\mathcal{V}_{0}^{d},\mathcal{V}_{1}^{d})]$ and a sequence of independent fair coin flips $\bm{e}=[{e}^{1},e^{2},\cdots,{e}^{d}]$, i.e. ${e}^{i}$ follows $\text{Bernoulli}(0.5)$, the multilayer reweighting is defined as:
@@
When vocabulary bipartitions $({\mathcal{V}_{0}}^{i},{\mathcal{V}_{1}}^{i})$ and fair coin flips $e_{i}$ are independent across layers, the multilayer reweighting is still unbiased, i.e., $\mathbb{E}_{\bm{e}}[P_{M,w}]=P_{M}$. A discussion and proof of this property can be found in App. A.3.
@@ 4.3 XOR-enhanced Position Allocation
Note that in this encoding process, message bits and one-time pad masks are contructed pseudorandomly for each token during text generation. The vocabulary partitions used in unbiased reweighting are constructed pseudorandomly before the entire text generation process. Compared to existing methods which construct vocabulary partitions or permutation for each token (Kirchenbauer et al., 2023; Hu et al., 2023; Wu et al., 2023), this approach is more efficient.
@@
$\hat{{e}}^{i}=\begin{cases}0&\text{if }x\in\mathcal{V}_{0}^{i},\\ 1&\text{if }x\in\mathcal{V}_{1}^{i}.\end{cases}$

### v2

@@ 1 Introduction
Large Language Models (LLMs) (OpenAI, 2022; Team, 2024b) have recently emerged in advancing cutting-edge technologies, and their rapid evolution has made LLM-generated content increasingly indistinguishable from human-created text. However, pressing security challenges (Zhang et al., 2023b; Wei et al., 2025; Gong et al., 2025; Luo et al., 2025; Zhang et al., 2025a, b, 2023a, 2024b; Pan et al., 2025) have been raised with the rapid progression of AI technologies (Zheng et al., 2024b; Zhang et al., 2024a; Zheng et al., 2024a, 2023, 2025; Feng & Dai, 2023), such as the misuse of LLMs (Tang et al., 2024) and synthetic data pollution (Shumailov et al., 2024). Therefore, recent legislation, such as the AI Act of the European Union (European Commission, 2023) and the executive order of the U.S. Department of Commerce (The White House, 2023), mandates the implementation of technical measures to mark and detect LLM-generated content. Thus, it appeals to LLM watermarking (Zhang et al., 2024d; Tang et al., 2024), a promising method to address these challenges, offering a proactive approach to embedding identifiable patterns in LLM-generated text (Liu et al., 2024).
@@
Inference-
Time
/ Aaronson (2023)
@@
Cumulative
Probability
$\gamma$-reweight (Hu et al., 2024)
DiPmark (Wu et al., 2024)
Token
Counting
GINSEW (Zhao et al., 2023)
@@
Depending on whether it is integrated with the language model, watermarking can be categorized into inference-time watermarking and post-hoc watermarking (Tang et al., 2024). Compared to post-hoc methods that require an additional auxiliary module (Abdelnabi & Fritz, 2021; Yang et al., 2022; Yoo et al., 2023), inference-time watermarking (Dathathri et al., 2024; Aaronson, 2023; Kirchenbauer et al., 2023) can embed detectable patterns linked to a secret key by modifying the LLM token sampling process in a plug-and-play manner, thereby efficiently achieving a cryptographically guaranteed detectable secret watermark (Aaronson, 2023; Kirchenbauer et al., 2023).
In general, implementing language model watermarking necessitates three key requirements: 1) Text quality preservation: The integration of watermarking mechanisms must maintain the fundamental utility and performance of language models (Hu et al., 2024; Kuditipudi et al., 2024). 2) Model-agnostic detection: The watermark verification process should function independently of the generating model’s architecture, parameters, or access rights (Kirchenbauer et al., 2023; Kuditipudi et al., 2024). 3) Message embedding capacity: The watermarking scheme should accommodate sufficient capacity to encode crucial metadata such as model identity, generation timestamp, and content provenance (Bob & Dan, 2024; Pang et al., 2024).
Note that it is not trivial to satisfy the above three requirements, as there is a trade-off between text quality preservation and message embedding capacity (Yoo et al., 2024; Qu et al., 2024; Kirchenbauer et al., 2023). Some existing works (Yoo et al., 2024; Qu et al., 2024) achieve efficient multi-bit watermarking by embedding one message bit per token through hashing. However, they rely on text quality-compromising zero-bit watermarking (Kirchenbauer et al., 2023) as a foundation, leading their enhancement of message embedding capacity inevitably come at the cost of text quality preservation. Note that, although other methods (Fernandez et al., 2023; Fairoze et al., 2025; Wang et al., 2024) can mitigate this issue to some extent by associating messages with secret keys or hash functions, they lack detection efficiency due to using messages non-invertibly.
@@ 2 Related Work
LLM watermarking aims to address the urgent need to differentiate model-generated text from human-written content. Text watermarking (Fang et al., 2017; Ziegler et al., 2019; He et al., 2022b; Fairoze et al., 2025; Fridrich et al., 2004; Christ et al., 2024; Zhao et al., 2025) has evolved significantly since its early application in content protection. Here we review related work on inference-time watermarking around the three key capacities in LLM watermarking.
@@
Text quality preservation. Early watermarking methods such as Soft Red List (Kirchenbauer et al., 2023) significantly influenced the field but faced challenges with text quality preservation due to their biased nature. Recent work has addressed this challenge through two main approaches: unbiased reweighting (Hu et al., 2024; Wu et al., 2024) and distortion-free sampling (Kuditipudi et al., 2024; Zhao et al., 2025; Dathathri et al., 2024). Unbiased reweighting methods, such as $\gamma$-reweight (Hu et al., 2024) and DiPmark (Wu et al., 2024), modify the probability distributions of generated text to inject watermarks while ensuring that the expected probability distribution remains unchanged. Distortion-free sampling methods, such as Gumbel sampling (Aaronson, 2023; Kuditipudi et al., 2024) and SynthID (Dathathri et al., 2024), employ secret keys to guide the sampling process rather than modifying the probability distribution. Many of them enable model-agnostic detection, but lack multi-bit functionality.
@@
Model-agnostic detection. The ability to detect watermarks without accessing the original or an auxiliary model is crucial for practical deployment. Existing watermarking methods developed from unbiased reweighting (Hu et al., 2024; Wu et al., 2024) utilize cumulative probability to partition the vocabulary, which makes them unsuitable for model-agnostic detection. In the scenario of multi-bit watermarking, a similar concept critical for practical deployment is message-agnostic detection. That is, the ability to extract messages without access to the message space. Among existing multi-bit watermarking methods, only the works of Yoo et al. (2024) and (Qu et al., 2024) satisfy this property through extracting messages from a voting matrix bit by bit. However, they are developed based on Soft Red List and thus without guaranteeing text quality.
@@
Message embedding capacity. Message embedding capacity enables information to be embedded in and extracted from watermarked text, which is essential for tracing text provenance (Cohen et al., 2025; Pang et al., 2024). Recent studies develop multi-bit watermarking for message embedding capacity by associating each message with a unique secret key (Fernandez et al., 2023), or incorporating messages into hash functions as inputs (Fairoze et al., 2025; Wang et al., 2024). These methods require access to message space for extraction, as the messages themselves are entangled with hash functions in a non-invertible manner.
@@
Yoo et al. (2024) proposed a position allocation technique incorporating Soft Red List (Kirchenbauer et al., 2023), which can solve this problem. They allocate each token to a subunit of a message for watermark embedding, and extract the message bit by bit, which is in a message-agnostic way.  Qu et al. (2024) further incorporates error correction codes to enhance multi-bit watermarking resilience. However, these methods were designed using Soft-List, leading to a lack of unbiasedness. As shown in the Tab. 1, a comprehensive comparison between BiMark and related studies is presented, highlighting our contributions in this field.
@@ 3.1 Watermarking for LLMs
Watermarking LLM-generated text. To reliably distinguish between LLM-generated and human-written text, watermarking methods actively inject detectable signals using secret keys (Kirchenbauer et al., 2023; Aaronson, 2023) into LLM-generated text. During detection, through measuring the similarity between extracted signals and injected signals with the same keys, the watermark can be verified. Early works (Kirchenbauer et al., 2023; Kuditipudi et al., 2024; Hu et al., 2024) only focus on addressing the binary question of whether a text contains watermarks, referred to zero-bit watermarking. For more reliable text authenticity, watermarking methods that can carry informative messages are proposed, referred to as multi-bit watermarking  (Yoo et al., 2024; Fairoze et al., 2025; Cohen et al., 2025).
@@ 3.2 Soft Red List Watermarking
If the z-score surpasses a predefined threshold, it suggests that the text is watermarked, since the proportion of green tokens deviates significantly from the expected value $\gamma$. When using the z-score for watermark detection, the corresponding p-value reflects the false positive rate (Type-I error), whereas the false negative rate (Type-II error) is influenced by the inherent entropy characteristics of the LLM (Kirchenbauer et al., 2023; Kuditipudi et al., 2024).
@@ 3.3 Watermarking via Unbiased Reweighting
Definition 3.1 ($n$-shot undetectable (Hu et al., 2024)).
@@
(Unbiased reweighting function (Hu et al., 2024)). Given a random variable $e$ and a reweighting function $R_{e}:\Delta_{\mathcal{V}}\rightarrow\Delta_{\mathcal{V}}$, we say that $R_{e}$ is an unbiased reweighting function if and only if for all probability distributions $P\in\Delta_{\mathcal{V}},~E_{e}[R_{e}(P)]=P$.
@@ 3.4 Multi-bit Watermarking via Position Allocation
Many prior works (Fernandez et al., 2023; Fairoze et al., 2025; Wang et al., 2024; Yoo et al., 2024; Qu et al., 2024) have explored multi-bit watermarking that extends beyond zero-bit detection. Among them, Yoo et al. (2024) proposed Multi-bit Watermarking via Position Allocation (MPAC), which enables message embedding without additional latency compared to Soft Red List and allows message extraction in a message-agnostic manner. We examine this technique, as it inspired BiMark, which integrates it with XOR operations to achieve unbiased watermarking.
@@ 4.1 Bit-Flip Unbiased Reweighting
An unbiased reweighting function is the core component of unbiased watermarking. While unbiased reweighting helps preserve text quality in watermarking, existing methods  (Hu et al., 2024; Wu et al., 2024) based on cumulative probability distributions require access to model probabilities, making them unsuitable for model-agnostic detection.
@@
Formally, let a random variable e{e} represent a fair coin flip, where e=1{e}=1 indicates heads and e=0{e}=0 indicates tails. Given a vocabulary bipartition $(\mathcal{V}_{0},\mathcal{V}_{1})$, the bit-flip reweighting function Fig. 2 adjusts an original probability as follows:
[…]
@@
For practical implementation, we introduce a base scaling factor $\tilde{\delta}\in[0,1]$ and derive $\delta_{1}$ adaptively based on $\tau$ from Eq. (3). However, two scenarios [^2] For ease of presentation, we assume $(1+\delta_{1})\tau>(1+\delta_{0})(1--\tau)$ and focus on the edge outlier caused by expanding $\mathcal{V}_{1}$. require special attention: (1) If $\tau=0$, we set $\delta_{1}=0$ since all tokens in $\mathcal{V}_{1}$ have zero probability and any non-zero $\delta_{1}$ would have no effect on the probability distribution and will break the symmetry of our reweighting scheme when e=1{e}=1. (2) If $(1+\tilde{\delta})\tau>1$, we set $\delta_{1}=(1-\tau)/\tau$ which makes the probability of $\mathcal{V}_{1}$ become $1$ and the probability of $\mathcal{V}_{0}$ become 0, ensuring the reweighted distribution remains a valid probability distribution over two partitions. Consequently, the scaling factor $\delta_{1}$ becomes:
[…]
@@
For any probability distribution over a vocabulary $\mathcal{V}$ and a reweighting function $R_{\theta,e}$ defined above, we have $E_{e}[R_{\theta,{e}}(P)]=P$.
@@ 4.2 Multilayer Unbiased Reweighting
Formally, given a sequence of independent vocabulary bipartitions $(\mathcal{V}_{0}^{1},\mathcal{V}_{1}^{1}),\cdots,(\mathcal{V}_{0}^{d},\mathcal{V}_{1}^{d})]$ and a sequence of independent fair coin flips $\bm{e}=[{e}^{1},e^{2},\cdots,{e}^{d}]$, i.e. ${e}^{i}$ follows $\text{Bernoulli}(0.5)$, the multilayer reweighting is defined as:
@@
When vocabulary bipartitions $({\mathcal{V}_{0}}^{i},{\mathcal{V}_{1}}^{i})$ and fair coin flips $e_{i}$ are independent across layers, the multilayer reweighting is still unbiased, i.e., $E_{\bm{e}}[P_{M,w}]=P_{M}$. A discussion and proof of this property can be found in App. A.3.
@@ 4.3 XOR-enhanced Position Allocation
Note that in this encoding process, message bits and one-time pad masks are contructed pseudorandomly for each token during text generation. The vocabulary partitions used in unbiased reweighting are constructed pseudorandomly before the entire text generation process. Compared to existing methods which construct vocabulary partitions or permutation for each token (Kirchenbauer et al., 2023; Hu et al., 2024; Wu et al., 2024), this approach is more efficient.
@@
[…]